Admitting New Partners Who Buy an Interest and Contribute Assets

The capital accounts of Trent Henry and Tim Chou have balances of $139,000 and $100,000, respectively. LeAnne Gilbert and Becky Clarke are to be admitted to the partnership. Gilbert buys one-fifth of Henry’s interest for $32,000 and one-fourth of Chou’s interest for $22,000. Clarke contributes $33,900 cash to the partnership, for which she is to receive an ownership equity of $33,900.

a1. Journalize the entry to record the admission of Gilbert. For a compound transaction, if an amount box does not require an entry, leave it blank.

Trent Henry, Capital
Tim Chou, Capital
LeAnne Gilbert, Capital

a2. Journalize the entry to record the admission of Clarke.

Cash
Becky Clarke, Capital

b. What are the capital balances of each partner after the admission of the new partners?

Partner Capital Balance
Trent Henry $
Tim Chou $
LeAnne Gilbert $
Becky Clarke $
Answer #1
a1
Trent Henry,Capital 27800 =139000*1/5
Tim Chou,Capital 25000 =100000*1/4
       LeAnne Gilbert,Capital 52800
a2
Cash 33900
       Becky Clarke,Capital 33900
b
Capital Balance
Trent Henry 111200 =139000-27800
Tim Chou 75000 =100000-25000
LeAnne Gilbert 52800
Becky Clarke 33900